OpenStudy (anonymous):

Solve the Equation: (dy/dx)= e^(x+y)/(y-1) I am positive this is a separable equation

OpenStudy (imstuck):

are you just trying to find the derivative of that function?

OpenStudy (freckles):

You know e^(x+y) can be written as e^x*e^y

OpenStudy (freckles):

\[\frac{dy}{dx}=e^x \cdot \frac{e^y}{y-1}\]

OpenStudy (freckles):

do you see how it is separable now? :)

OpenStudy (anonymous):

I tried that and then I had issues integrating it. The dy part not the dx, I can do dx. Would I use integration by parts for dy?

OpenStudy (freckles):

yes

OpenStudy (freckles):

You know you can write (y-1)/e^y as (y-1)*e^{-y}?

OpenStudy (anonymous):

alrighty I will try that and yes I know :)

OpenStudy (anonymous):

I got (-y/e^y)-e^x=c

OpenStudy (freckles):

that looks good

OpenStudy (freckles):

you can also write it without all of negatives y/e^y+e^x=-c or y/e^y+e^x=C <--

OpenStudy (anonymous):

alrighty thank you!

OpenStudy (freckles):

np